# An experimental platform for levitated mechanics in space

### v2

@@ Abstract
Conducting levitated mechanical experiments in extreme conditions has long been the aim of researchers, as it allows for the investigation of new fundamental physics phenomena. One of the great frontiers has been sending these experiments into the micro-g environment of space, with multiple proposals calling for such a platform. At the same time, levitated sensors have demonstrated a high sensitivity to external stimuli which will only improve in low-vibrational conditions. conditions This paper describes the development of a technology demonstrator for optical and magnetic trapping experiments in space. Our payload represents the first concrete step towards future missions with aims of probing fundamental physical questions: matter-wave interferometry of nanoparticles to probe the limits of macroscopic quantum mechanics, detection of Dark Matter candidates and gravitational waves to test physics beyond the Standard Model, and accelerometry for Earth-observation.
@@ 1 Introduction
  • Longer coherence and free evolution times are required for observing the quantum properties of a released nanoparticle’s spatial evolution. The trapping laser’s interaction with the nanoparticle acts a measurement, causing wave-function collapse such that nanoparticles must be released from the trap to allow them to freely evolve. As the required free evolution time for larger nanoparticles ranges between 100 ms to 100 s [26], terrestrial experiments cannot conduct these tests as gravity would accelerate the particles out of the trapping region during this period. In space, this issue is removed, such that experiments are limited solely by the control of the nanoparticle’s initial thermal motion [27].
@@ 4 Payload Design
One of the payload’s end-plates was designed as a heat sync for the laser driver PCB and FPGA (see Section 5), while the other was recessed to hold the optical trap’s photodetector (see Section 6). Both end-plates were bolted to the stainless steel baseplate which had was designed with M10 through-holes for bolting the payload to the Nyx capsule. The end-plates were also held together at the top of the payload by aluminium stringers, while larger stringers between the two end-plates held the vacuum chamber near the payload’s midpoint. Another stringer was added between the top stringers to hold the optical fibre mating sleeves. Stainless steel cover-plates were attached to the stringers and end-plates to enclose the payload. Images of the completed payload can be seen in Fig. 3. To ensure that all structural components would be secure throughout the mission, all threaded parts used locking threads (Spiralock) and a threadlocking agent (Loctite 242). Similarly, all internal wires and optical fibres were secured with dots of epoxy (3M DP190) at least every 2 cm.
@@ 5 Electronics
The microcontroller has 3 internal analogue-digital converters which are used to collect data from the two photodetectors and an reference accelerometer (Analogue Devices EVAL-ADXL354BZ). The data will be collected with 16-bit resolution in 8 MB packets (limited by the microcontroller’s RAM) and saved with duplicates on an internal micro-SD card to protect the data from corruption. Before reaching the microcontroller, the photodetectors’ signals are passed through band-pass filters, DC-shifted and amplified to maximize the system’s sensitivity. The data from the optical and magnetic traps in acquired, respectively, at sampling rates of 500 kSa/s and 10 kSa/s to allow for sufficient bandwidth to measure the predicted oscillation frequencies of each trap. The accelerometer data is only saved alongside the magnetic trap data, as the nanoparticle is too light to effectively measure the same vibrational noise floor.
@@ 7 Proof of Functionality
While the payload was undergoing software testing (see Section 8), data was taken with the magnetic trap in the payload using all the systems that will be used throughout the mission. The PSD of the graphite’s motion (shown in Fig. 4 shows that the graphite’s multiple distinct oscillation frequencies can be detected and match well with the predicted frequency range of 1-100 Hz.
@@ Acknowledgments
We thank James Chalk, Philip Connell, Mark Bampton, Damon Grimsey, Jamie Faux, Oliver Warner, Leigh Allen, Luca Ferrain, Colin Chau, and Robin Elliot for their expert technical support and consultation. We also thank the Space South Central cluster for their support. This project has been made possible by the ESA Payload Master’s competition, providing us with a place in The Exploration Company’s ‘Mission Possible’ launch. HU acknowledges funding by the EU EIC Pathfinder project QuCoM (101046973). We thank for support the UKRI EPSRC (EP/W007444/1, EP/V000624/1 and EP/X009491/1), the Leverhulme Trust (RPG-2022-57) and the QuantERA II Programme (project LEMAQUME) that has received funding from the European Union’s Horizon 2020 research and innovation programme under Grant Agreement No 101017733. CV acknowledges support from DLR project NaiS (50WM2180).

### v3

@@ Abstract
Conducting levitated mechanical experiments in extreme conditions has long been the aim of researchers, as it allows for the investigation of new fundamental physics phenomena. One of the great frontiers has been sending these experiments into the micro-g environment of space, with multiple proposals calling for such a platform. At the same time, levitated sensors have demonstrated a high sensitivity to external stimuli which will only improve in low-vibrational conditions. This paper describes the development of a technology demonstrator for optical and magnetic trapping experiments in space. Our payload represents the first concrete step towards future missions with aims of probing fundamental physical questions: matter-wave interferometry of nanoparticles to probe the limits of macroscopic quantum mechanics, detection of Dark Matter candidates and gravitational waves to test physics beyond the Standard Model, and accelerometry for Earth-observation.
@@ 1 Introduction
  • Longer coherence and free evolution times are required for observing the quantum properties of a released nanoparticle’s spatial evolution. The trapping laser’s interaction with the nanoparticle acts as a measurement, causing wave-function collapse such that nanoparticles must be released from the trap to allow them to freely evolve. As the required free evolution time for larger nanoparticles ranges between 100 ms to 100 s [26], terrestrial experiments cannot conduct these tests as gravity would accelerate the particles out of the trapping region during this period. In space, this issue is removed, such that experiments are limited solely by the control of the nanoparticle’s initial thermal motion [27].
@@ 4 Payload Design
One of the payload’s end-plates was designed as a heat sync for the laser driver PCB and FPGA (see Section 5), while the other was recessed to hold the optical trap’s photodetector (see Section 6). Both end-plates were bolted to the stainless steel baseplate which was designed with M10 through-holes for bolting the payload to the Nyx capsule. The end-plates were also held together at the top of the payload by aluminium stringers, while larger stringers between the two end-plates held the vacuum chamber near the payload’s midpoint. Another stringer was added between the top stringers to hold the optical fibre mating sleeves. Stainless steel cover-plates were attached to the stringers and end-plates to enclose the payload. Images of the completed payload can be seen in Fig. 3. To ensure that all structural components would be secure throughout the mission, all threaded parts used locking threads (Spiralock) and a threadlocking agent (Loctite 242). Similarly, all internal wires and optical fibres were secured with dots of epoxy (3M DP190) at least every 2 cm.
@@ 5 Electronics
The microcontroller has 3 internal analogue-digital converters which are used to collect data from the two photodetectors and a reference accelerometer (Analogue Devices EVAL-ADXL354BZ). The data will be collected with 16-bit resolution in 8 MB packets (limited by the microcontroller’s RAM) and saved with duplicates on an internal micro-SD card to protect the data from corruption. Before reaching the microcontroller, the photodetectors’ signals are passed through band-pass filters, DC-shifted and amplified to maximize the system’s sensitivity. The data from the optical and magnetic traps is acquired, respectively, at sampling rates of 500 kSa/s and 10 kSa/s to allow for sufficient bandwidth to measure the predicted oscillation frequencies of each trap. The accelerometer data is only saved alongside the magnetic trap data, as the nanoparticle is too light to effectively measure the same vibrational noise floor.
@@ 7 Proof of Functionality
While the payload was undergoing software testing (see Section 8), data was taken with the magnetic trap in the payload using all the systems that will be used throughout the mission. The PSD of the graphite’s motion (shown in Fig. 4) shows that the graphite’s multiple distinct oscillation frequencies can be detected and match well with the predicted frequency range of 1-100 Hz.
@@ Acknowledgments
We thank James Chalk, Philip Connell, Mark Bampton, Damon Grimsey, Jamie Faux, Oliver Warner, Leigh Allen, Luca Ferrain, Colin Chau, and Robin Elliott for their expert technical support and consultation. We also thank the Space South Central cluster for their support. This project has been made possible by the ESA Payload Master’s competition, providing us with a place in The Exploration Company’s ‘Mission Possible’ launch. HU acknowledges funding by the EU EIC Pathfinder project QuCoM (101046973). We thank for support the UKRI EPSRC (EP/W007444/1, EP/V000624/1 and EP/X009491/1), the Leverhulme Trust (RPG-2022-57) and the QuantERA II Programme (project LEMAQUME) that has received funding from the European Union’s Horizon 2020 research and innovation programme under Grant Agreement No 101017733. CV acknowledges support from DLR project NaiS (50WM2180).